# EmoSteer-TTS: Fine-Grained and Training-Free Emotion-Controllable Text-to-Speech via Activation Steering

### v2

@@ Introduction
Datasets. To obtain effective steering vectors, we construct a curated emotional speech dataset by collecting samples with clear emotional expression from multiple corpora: MSP-Podcast (Lotfian and Busso 2017), IEMOCAP (Busso et al. 2008), RAVDESS (Livingstone and Russo 2018), CREMA-D (Cao et al. 2014), TESS (Pichora-Fuller and Dupuis 2020), SAVEE (Jackson and Haq 2014), ASVP-ESD (Landry et al. 2020), CASIA (CASIA 2023), M3ED (Zhao et al. 2022), ESD (Zhou et al. 2022), and Emo-Emilia (Zhao et al. 2025). The resulting dataset contains 6,900 utterances covering six basic emotions (anger, happiness, sadness, disgust, surprise, fear) and neutrality. Each emotion includes 1,000 samples, 500 in English and 500 in Chinese, except for fear, which has 400. The dataset includes diverse speakers with balanced gender distribution. The construction details are provided in Appendix B of the Supplementary Materials. This dataset is used to compute activation differences between neutral and emotional speech, as defined in Eq. 1. To identify the top-$k$ tokens for each emotion, we synthesize speech using 10 neutral ESD samples.
@@
Models. We enhance three SOTA flow matching-based TTS models (F5-TTS, CosyVoice2, E2-TTS (Eskimez et al. 2024)) using our proposed method, and compare their controllability with that of leading EC-TTS baselines, including both label-based methods with adjustable control strength (EmoSphere++, EmoDubber, HED-TTS (Inoue et al. 2025)) and description-based methods (EmoVoice, CosyVoice2, FleSpeech (Li et al. 2025)). Appendix C provides detailed model and hardware configurations for all experiments.
Emotion Conversion and Interpolation
@@
Emotion Conversion. We conduct emotion conversion using 100 neutral reference speech samples (50 English from MSP-Podcast and 50 Chinese from ESD), with $\alpha$=2.0 and kk=200. We report Word Error Rate (WER), Speaker Similarity (S-SIM), Emotion Similarity (E-SIM), and Naturalness Mean Opinion Score (N-MOS, 1–5 scale, rated by 8 participants, see Appendix D.1 for details). WER is derived from Whisper-Large V3 (Radford et al. 2023) transcriptions. S-SIM is the cosine similarity between embeddings of synthesized and neutral reference from an SER model (Bredin et al. 2020). E-SIM is computed as the cosine similarity between emotion2vec embeddings of synthesized speech and 100 anchor emotional samples (per emotion) from MSP-Podcast and ESD. We compute baseline scores using their demo samples for fairness. As shown in Table 1, EmoSteer-TTS achieves superior performance across multiple methods. Integrated with F5-TTS, it yields a low WER of 2.79, close to CosyVoice2 (2.53) and far better than label-based baselines. It also maintains high S-SIM (0.64, 0.65), indicating strong speaker preservation. F5-TTS and E2-TTS with EmoSteer-TTS reach the top E-SIM scores (0.29, 0.28), outperforming all baselines and matching FleSpeech. In N-MOS, “CosyVoice2 + EmoSteer-TTS” (3.65) is close to the best (EmoVoice, 3.81), and our method consistently outperforms label-based systems. Fig. 4(a) also shows the shift in emotion2vec probability distribution (averaged across three models) for 100 synthesized samples per emotional tone before ($\alpha$=0) and after ($\alpha$=2) emotion conversion.
@@
Emotion Interpolation. We reuse the speech samples from the emotion conversion experiments to perform interpolation (kk=200), gradually shifting emotional tone from neutrality to a target emotion. To assess fine-grained controllability, we report the Emotion Interpolation MOS (EI-MOS; 1–5 scale, 8 participants), which evaluates the alignment between target intensity and synthesized speech. Detailed criteria for EI-MOS are provided in Appendix D.2. Label-based baselines use intensity levels (e.g., 0.5 or 1.0) to control, while description-based methods, lacking intensity control, are excluded in this experiment. For fairness, baseline metrics are computed using their official demo samples. As shown in Table 1, EmoSteer-TTS achieves higher EI-MOS than label-based baselines, indicating superior capability in controlling emotional intensity. Notably, “F5-TTS+EmoSteer-TTS” obtains the highest EI-MOS of 4.00, outperforming EmoSphere++ and HED-TTS, showing better alignment with intended emotion levels. E2-TTS and CosyVoice2 variants also perform well, suggesting EmoSteer-TTS generalizes across different models. As shown in Fig. 4(b), the average predicted emotion probabilities (via emotion2vec) vary smoothly with $\alpha$, illustrating EmoSteer-TTS’s fine-grained controllability. However, large $\alpha$ values (e.g., 3) may lead to unintelligible speech. Fig. 5(a) also illustrates smooth F0 transitions with increasing anger intensity. More examples are in Appendix E.1.
@@
Figure 5: Visualization of F0 contours. (a) An example showing how the F0 contour varies with steering intensity; (b) The speech tone becomes calmer after emotion erasure.
Emotion Erasure
@@
We randomly select 100 unseen emotional speech samples for each type of emotion from MSP-Podcast and ESD, and erase the emotional tone using Eq. 9. We ask 8 participants to rate the synthesized speech by giving an Emotion-Erasure MOS (EE-MOS, 1-5 scale), which indicates how well the synthesized speech reflects the intended emotion erasure. Higher EE-MOS reflects better erasure performance. The standard for EE-MOS is detailed in Appendix D.3. We set $\beta$=2.5 and kk=200 for this experiment. As shown in Table 1, our method achieves a relatively high EE-MOS score, indicating effective removal of the target emotion. The decreased target emotion scores predicted by emotion2vec in Fig. 4(c) further demonstrate the emotion erasing ability of EmoSteer-TTS. Fig. 5(b) illustrates the variation of F0 contours when gradually erasing an emotional tone. More visualizations are provided in Appendix E.2.
Composite Control
@@
Multi-Emotion Steering. We use the same neutral samples from the emotion conversion experiment as reference speech. For simplicity, this experiment simultaneously adds two emotions to the synthesized speech ($\alpha_{1}$$\alpha_{2}$=2, kk=200). As shown in Fig. 4(e), the predicted emotion distributions align closely with the intended emotion pairs. For example, the row labeled “F, H” shows elevated probabilities for both fear (0.22) and happiness (0.33), while “Sa ,Su” leads to strong activations for sadness (0.28) and surprise (0.42). These results indicate that EmoSteer-TTS can effectively blend multiple emotions, enabling expressive and nuanced speech synthesis beyond single-label control.
Analysis of Emotion Steering Dynamics
@@
In this subsection, we analyze the emotion steering dynamics of our method. All analyses are conducted on F5-TTS, which consists of 22 DiT block layers and performs 32 flow matching steps to generate the final mel-spectrograms.
@@
The Impact of Top-$k$. The parameter $k$ determines the number of emotion-related tokens selected to construct the steering vector $\mathbf{s}^{l}$ at layer $l$. A larger $k$ introduces more tokens into the steering signal, potentially capturing a broader range of emotional nuances, while a smaller $k$ focuses on the most dominant emotion features. We conduct emotion conversion with varying $k$ values (e.g., $k\in{10,50,...,400}$) and evaluate the impact on the emotion of synthesized speech. As shown in Fig. 4(f), increasing $k$ generally leads to higher average emotion probabilities across all categories, particularly for anger and happiness, which peak around kk=200. This indicates that incorporating more emotion-relevant tokens enriches the steering signal and enhances emotional expressiveness. However, gains diminish beyond kk=200, suggesting a saturation point where additional tokens offer limited benefit. Therefore, we set kk=200 in all main experiments to balance effectiveness and efficiency.
Steering Different Layers. We investigate how the choice of controlled layer influences the effectiveness of emotion conversion. We apply the steering vector $\mathbf{s}^{l}$ at shallow layers (1-7), middle layers (8-14), deep layers (15-22), or multiple spaced layers (1, 6, 11, 16, 21) and compare the emotion of synthesized speech. As shown in Fig. 4(g), applying steering vectors at shallow layers yields moderate average emotion probabilities, indicating an initial but limited emotional influence. At middle layers, these probabilities increase slightly, suggesting a stronger yet still evolving emotional control. Interestingly, deep layers show a slight decline, implying they may focus on fine-grained emotional nuances rather than amplifying emotion intensity. The most notable effect arises when steering is applied to multiple spaced layers, resulting in a clear and significant boost in emotional probabilities across six emotions. This suggests a cumulative, distributed influence, enabling more expressive synthesis. In summary, emotion steering varies with depth: shallow to deep layers offer progressively refined but localized control, while a multi-layered strategy produces the most robust and holistic emotional expression.
Steering Different Flow Matching Steps. F5-TTS involves 32 conditional flow matching (CFM) steps that gradually refine the mel-spectrogram prediction. To understand how steering at different stages affects emotion synthesis, we apply emotion control to varying subsets of the steps (e.g., early: steps 0–10; middle: steps 11–21; late: steps 22–32; or all steps). As shown in Fig. 4(h), emotion steering during early steps yields minimal effect, with low average emotion probabilities. In contrast, middle and late-stage steering leads to a marked increase, indicating greater influence as the spectrogram solidifies. The strongest emotional expression arises when steering spans all steps, highlighting the importance of continuous guidance throughout the refinement process. This aligns with the design of the CFM model, which conditions on reference speech at every step, necessitating activation modulation across the entire flow to achieve effective emotion control. Therefore, we apply emotion steering across all steps in the main experiments, i.e., 32 steps for F5-TTS and E2-TTS, and 10 steps for CosyVoice2.
Discussion
By steering models’ activations across layers and generation steps, our method offers fine-grained modulation of emotional tone without retraining. This design makes the method broadly applicable to various flow matching-based TTS models. However, its effectiveness depends on the quality of the datasets used to compute steering vectors. Steering across multiple layers or steps also increases computational cost, and handling subtle emotions remains a challenge. Future work will focus on enhancing the emotion representation space using multimodal or self-supervised learning and investigating adaptive steering strategies that dynamically allocate control across layers and steps based on input content and target speaking styles. In addition, the proposed method can also be utilized to construct speech datasets with fine-grained emotion annotations, which is crucial for developing more expressive TTS methods.
Conclusion
@@
We presented EmoSteer-TTS, a training-free framework for fine-grained, continuous, and interpretable emotion control in speech synthesis. By steering a subset of internal activations in a TTS model, our method enables flexible emotional manipulation, including emotion conversion, interpolation, and erasure, without modifying or fine-tuning the pretrained TTS model. We also constructed a curated emotional speech dataset to support steering vector construction. Extensive experiments confirm that EmoSteer-TTS achieves robust, zero-shot emotion control with broad applicability, outperforming SOTA methods. The analysis also offers deeper insights into the emotion steering dynamics of flow matching-based TTS. To the best of our knowledge, this is the first fine-grained EC-TTS approach that needs no additional training.
Appendix A: Main Code for Fine-Grained Emotion Control

### v3

@@ 4.1 Datasets and Models
Datasets for Steering Vector Construction. To obtain effective steering vectors, we construct a curated emotional speech dataset by collecting samples with clear emotional expression from multiple corpora: MSP-Podcast (Lotfian & Busso, 2017), IEMOCAP (Busso et al., 2008), RAVDESS (Livingstone & Russo, 2018), CREMA-D (Cao et al., 2014), TESS (Pichora-Fuller & Dupuis, 2020), SAVEE (Jackson & Haq, 2014), ASVP-ESD (Landry et al., 2020), CASIA (CASIA, 2023), M3ED (Zhao et al., 2022), ESD (Zhou et al., 2022), and Emo-Emilia (Zhao et al., 2025). The resulting dataset contains 6,900 utterances covering six basic emotions (anger, happiness, sadness, disgust, surprise, fear) and neutrality. Each emotion includes 1,000 samples, 500 in English and 500 in Chinese, except for fear, which has 400. The dataset includes diverse speakers with a balanced gender distribution. The construction details are provided in Appendix B. This dataset is used to compute activation differences between neutral and emotional speech, as defined in Eq. 1. To identify the top-$k$ tokens for each emotion, we synthesize speech using 10 random neutral ESD samples as references (5 English and 5 Chinese).
Datasets for Inference-Time Emotion Steering. 1) In-distribution evaluation: We sample neutral and emotional reference speeches from MSP-Podcast and ESD, which are excluded from steering vector computation. 2) Out-of-distribution (OOD) evaluation: We sample neutral speech from SeedTTS Anastassiou et al. (2024) test sets and emotional speech from EMNS Noriy et al. (2023).
@@
Models. We enhance three SOTA flow matching-based TTS models (F5-TTS, CosyVoice2, E2-TTS (Eskimez et al., 2024)) using our proposed method, and compare their controllability with that of leading EC-TTS baselines, including both label-based methods with adjustable control strength (EmoSphere++, EmoDubber, HED-TTS (Inoue et al., 2025)) and description-based methods (EmoVoice, CosyVoice2, FleSpeech (Li et al., 2025)). Appendix C provides detailed model and hardware configurations for all experiments.
4.2 Emotion Conversion and Interpolation
Table 1: In-distribution and OOD comparison with emotion-controllable baselines.
Method Conversion ($\alpha=2.0$) Interpolation Erasure ($\beta=2.5$)
WER($\downarrow$) S-SIM($\uparrow$) E-SIM($\uparrow$) N-MOS($\uparrow$) EI-MOS($\uparrow$) E-SIM($\uparrow$) $E$-MOS($\uparrow$)
emotion2vec / SenseVoice emotion2vec / SenseVoice
Label-based*
EmoSphere++ 16.25 0.44 0.25 / 0.24avg=0.245 $3.23_{\pm 0.81}$ 3.50±1.05 - -
EmoDubber 18.61 0.41 0.25 / 0.22avg=0.235 2.47±1.22 2.21±1.08 - -
HED-TTS 13.27 0.52 0.22 / 0.26avg=0.240 3.31±0.79 2.59±0.76 - -
Description
-based*
EmoVoice 2.91 0.58 0.27 / 0.25avg=0.260 3.81±0.86 - - -
CosyVoice2 2.53 0.73 0.24 / 0.27avg=0.255 3.69±1.07 - - -
FleSpeech 9.34 0.54 0.29 / 0.26avg=0.275 3.07±0.75 - - -
Unsteered F5-TTS 2.14 0.66 0.07 / 0.04avg=0.055 3.79±0.89 - 0.03 / 0.05avg=0.040 1.21±1.17
E2-TTS 2.71 0.64 0.05 / 0.08avg=0.065 3.51±0.94 - 0.06 / 0.02avg=0.040 1.35±1.05
In-distribution evaluation on MSP-Podcast (25% en) and ESD (25% en, 50% zh)
EmoSteer-TTS#
+ F5-TTS 2.79 0.64 0.29 / 0.26avg=0.275 3.29±1.05 4.00±0.89 0.27 / 0.25avg=0.260 4.02±0.85
+ E2-TTS 3.28 0.59 0.28 / 0.28avg=0.280 3.31±0.97 3.38±1.09 0.24 / 0.26avg=0.250 3.63±1.17
+ CosyVoice2 2.83 0.65 0.26 / 0.29avg=0.275 3.65±1.08 3.56±1.15 0.26 / 0.25avg=0.255 3.94±0.97
Cross-datasets (OOD) evaluation on EMNS (25% en) and SeedTT test sets (25% en, 50% zh)
EmoSteer-TTS#
+ F5-TTS 2.65 0.65 0.25 / 0.27avg=0.260 3.58±1.04 3.46±1.08 0.25 / 0.22avg=0.235 3.92±0.99
+ E2-TTS 3.41 0.55 0.26 / 0.25avg=0.255 3.44±1.07 3.50±0.97 0.24 / 0.27avg=0.255 3.57±1.03
+ CosyVoice2 2.86 0.66 0.28 / 0.25avg=0.265 3.49±1.01 3.48±1.27 0.23 / 0.21avg=0.220 3.98±0.94
*: Training-based, #: Training-free, -: Neither label-based, description-based, nor unsteered methods support interpolation or erasure.
The top three results are indicated in boldface. Unsteered backbones are shown in gray for reference.
@@ 4.2 Emotion Conversion and Interpolation
Emotion Conversion. We conduct emotion conversion using 100 neutral reference speech samples (50 English from MSP-Podcast and 50 Chinese from ESD), with $\alpha$=2.0 and kk=200. We report Word Error Rate (WER), Speaker Similarity (S-SIM), Emotion Similarity (E-SIM), and Naturalness Mean Opinion Score (N-MOS, 1–5 scale, see Appendix D for details). WER is derived from Whisper-Large V3 (Radford et al., 2023) transcriptions. S-SIM is the cosine similarity between the embeddings of synthesized and neutral reference from a speaker embedding model (Bredin et al., 2020). E-SIM is computed as the cosine similarity between emotion2vec embeddings of synthesized speech and 100 anchor emotional samples (per emotion) from MSP-Podcast and ESD. To mitigate potential metric overfitting from emotion2vec, we also report E-SIM scores computed with SenseVoice An et al. (2024) embeddings. Since we cannot guarantee the synthesis quality of reproduced baselines, we compute their scores using demo samples for fairness. The reproduced baseline results are additionally reported in Appendix E. As shown in Table 1, EmoSteer-TTS achieves superior performance across multiple methods. Integrated with F5-TTS, it yields a low WER of 2.79, close to CosyVoice2 (2.53) and far better than label-based baselines. It also maintains high S-SIMs (0.66, 0.65), indicating strong speaker preservation. F5-TTS, E2-TTS, and CosyVoice2 with EmoSteer-TTS reach the top E-SIM scores, outperforming all baselines and matching FleSpeech. In N-MOS, “EmoSteer-TTS+CosyVoice2” (3.65) is close to the best (EmoVoice, 3.81), and our method consistently outperforms label-based systems. Fig. 4(a) also shows the shift in emotion probability distribution (averaged across three models) for 100 synthesized samples per emotional tone before ($\alpha$=0) and after ($\alpha$=2) emotion conversion.
Figure 4: Emotion steering results on MSP-Podcast and ESD. ⚫ emotion2vec, ▲ SenseVoice.
@@
Emotion Interpolation. We reuse the speech samples from the emotion conversion experiments to perform interpolation (kk=200), gradually shifting emotional tone from neutrality to a target emotion. To assess fine-grained controllability, we report the Emotion Interpolation MOS (EI-MOS; 1–5 scale), which evaluates the alignment between target intensity and synthesized speech. Detailed criteria for EI-MOS are provided in Appendix D. Label-based baselines use intensity levels (e.g., 0.5 or 1.0) to control, while description-based methods, lacking intensity control, are excluded in this experiment. For fairness, baseline metrics are computed using their official demo samples. As shown in Table 1, EmoSteer-TTS achieves higher EI-MOS than label-based baselines, indicating superior capability in controlling emotional intensity. Notably, “EmoSteer-TTS+F5-TTS” obtains the highest EI-MOS of 4.00, outperforming EmoSphere++ and HED-TTS, showing better alignment with intended emotion levels. E2-TTS and CosyVoice2 variants also perform well, suggesting EmoSteer-TTS generalizes across different models. As shown in Fig. 4(b), the average predicted emotion probabilities (via emotion2vec and SenseVoice) vary smoothly with $\alpha$, illustrating EmoSteer-TTS’s fine-grained controllability. However, we find that large $\alpha$ values (e.g., 3) may lead to unintelligible speech. Fig. 5(a) also illustrates smooth F0 transitions with increasing anger intensity. More examples are provided in Appendix F.
4.3 Emotion Erasure
We randomly select 100 unseen emotional speech samples for each type of emotion from MSP-Podcast (50 English) and ESD (50 Chinese), and erase the emotional tone using Eq. 9. We report the average E-SIM between the emotionally erased samples and 100 randomly selected neutral samples from MSP-Podcast (50 English) and ESD (50 Chinese). We also report Emotion-Erasure MOS (EE-MOS, 1-5 scale), which indicates how well the synthesized speech reflects the intended emotion erasure. Higher EE-MOS reflects better erasure performance. The standard for EE-MOS is detailed
@@ 4.3 Emotion Erasure
Figure 5: Visualization of F0 contours. (a) An example showing how the F0 contour varies with steering intensity; (b) The speech tone (F0 contour) becomes calmer after emotion erasure.
Figure 6: Results of composite control: (a) emotion replacement and (b) multi-emotion steering (Abbreviations: Anger, Disgust, Fear, Happiness, Sadness, Surprise)
@@
in Appendix D. We set $\beta$=2.5, kk=200 for this experiment. As shown in Table 1, our method achieves a fairly high EE-MOS score, indicating effective removal of target emotions. The decreased target emotion scores shown in Fig. 4(c) further demonstrate the emotion erasing ability. Fig. 5(b) illustrates the variation of F0 contours when gradually erasing an emotional tone. Appendix F provides more visualizations.
4.4 Composite Control
@@
Multi-Emotion Steering. We use the same neutral samples from the emotion conversion experiment as reference speech. For simplicity, this experiment simultaneously adds two emotions to the synthesized speech ($\alpha_{1}$$\alpha_{2}$=2, kk=200). As shown in Fig. 6(b), the predicted emotion2vec distributions align closely with the intended emotion pairs. For example, the row labeled “F, H” shows elevated probabilities for both fear (0.22) and happiness (0.33), while “Sa ,Su” leads to strong activations for sadness (0.28) and surprise (0.42). These results indicate that EmoSteer-TTS can blend multiple emotions, enabling expressive and nuanced speech synthesis beyond single-label control.
4.5 Cross-Datasets Evaluation
Since some samples used for computing steering vectors come from the same datasets (e.g., MSP-Podcast, ESD), we also evaluate EmoSteer-TTS in an OOD setting. For emotion conversion and interpolation, we sample 100 neutral utterances from SeedTTS and 100 emotional anchors per emotion from EMNS; for emotion erasure, we use 100 emotional utterances from EMNS as references and 100 neutral anchors per emotion from SeedTTS. As shown in Table 1 (lower section), EmoSteer-TTS maintains robust performance on unseen datasets, with minimal degradation across metrics, demonstrating strong generalization beyond the steering data.
4.6 Analysis of Emotion Steering Dynamics
Figure 7: Analysis of emotion steering dynamics using emotion2vec predictions.
@@ 4.6 Analysis of Emotion Steering Dynamics
In this subsection, we analyze the emotion steering dynamics of our method. All analyses are conducted on F5-TTS, which consists of 22 DiT block layers and performs 32 flow matching steps to generate mel-spectrograms. We use the same neutral samples from the emotion conversion experiment as reference speech. We report emotion2vec emotion probabilities for all analyses.
@@
The Impact of Top-$k$. The parameter $k$ determines the number of emotion-related tokens used to construct the steering vectors. A larger $k$ introduces more tokens into the steering signal, potentially capturing a broader range of emotional nuances, while a smaller $k$ focuses on the most dominant emotion features. We conduct emotion conversion with varying $k$ values (e.g., $k\in{10,50,...,400}$) and evaluate their impact on the synthesized emotion. Fig. 7(a) shows that increasing $k$ generally leads to higher average emotion probabilities across all categories, particularly for anger and happiness, which peak around kk=200. Incorporating more emotion-relevant tokens enriches the steering signal, but gains plateau beyond $k=200$ for most emotions. We therefore use $k=200$ in all main experiments to balance expressiveness and efficiency.
Steering Different Layers. We examine how controlled layers affect emotion conversion by applying steering vectors $\mathbf{s}^{l}$ at shallow (1–7), middle (8–14), deep (15–22), and spaced layers (1, 6, 11, 16, 21). As shown in Fig. 7(b), shallow layers yield moderate emotional influence, middle layers provide slightly stronger control, and deep layers show a decline, likely focusing on acoustic details rather than emotion. Steering multiple spaced layers, however, significantly boosts probabilities across all six emotions. Overall, shallow-to-deep layers provide progressively refined control, and multi-layer steering enables the most effective emotion modulation.
Steering Different Flow Matching Steps. F5-TTS generates mel-spectrograms through 32 conditional flow-matching (CFM) steps. To assess the impact of steering at different stages, we apply emotion control to early (0–10), middle (11–21), late (22–32), or all steps. As shown in Fig. 7(c), early steering has little effect, while middle and late stages exert stronger influence as the spectrogram takes shape. The strongest emotion emerges when steering spans all steps, consistent with CFM’s stepwise conditioning on reference speech. Therefore, we apply emotion steering across all steps in the main experiments: 32 for F5-TTS and E2-TTS, and 10 for CosyVoice2.
5 Conclusion
@@
We presented EmoSteer-TTS, the first training-free framework for fine-grained, continuous, and interpretable emotion control in speech synthesis. By steering a subset of internal activations in a TTS model, our method enables flexible emotional manipulation, including emotion conversion, interpolation, and erasure, without modifying or fine-tuning the pretrained TTS model. We also constructed a curated emotional speech dataset to support steering vector construction. Extensive experiments confirm that EmoSteer-TTS achieves robust, zero-shot emotion control with broad applicability, outperforming SOTA methods. The analysis also offers deeper insights into the emotion steering dynamics of flow matching-based TTS. To the best of our knowledge, this is the first fine-grained EC-TTS approach that needs no additional training.
Limitations and Future Work. A limitation of our method is the reliance on high-quality emotional speech samples, albeit in modest quantities, to extract effective steering vectors. In addition, strong activation steering may introduce artifacts. Future work will explore combining activation steering with learning-based approaches to mitigate these issues.
6 Reproducibility Statement
To ensure the reproducibility of our work, we have provided comprehensive details throughout the paper and its appendices. Our proposed methodology, EmoSteer-TTS, is thoroughly described in Section 4, including the key algorithms for activation extraction, steering vector construction, and fine-grained control, accompanied by precise mathematical formulations. Appendix A further offers detailed code snippets illustrating the implementation of our core steering operations. Details regarding the datasets used for both steering vector construction and evaluation are presented in Section 4.1, with the specific curation and filtering process for our emotional speech dataset outlined in Appendix B. We also provide the code for dataset preprocessing and the processed dataset in the Supplementary Materials. The configurations for the TTS models (F5-TTS, E2-TTS, Cosy Voice2), including the specific layers and steps selected for steering, are detailed in Appendix C. The hyperparameters and experimental settings for all evaluations are specified within the relevant subsections of Section 4, and the criteria for our subjective evaluation metrics are defined in Appendix D. We will release the fully runnable code and curated dataset upon the paper’s acceptance to facilitate further research.
Appendix A Appendix A: Code Snippets for Fine-Grained Emotion Control